OpenStudy (alexh107):

Use mathematical induction to prove the statement is true for all positive integers n. The integer n^3 + 2n is divisible by 3 for every positive integer n.

OpenStudy (alexh107):

So far I have: For n = 1 1^3 + 2(1) / 3 = 3/3 = 1 True For n = k k^3 + 2k / 3 n = k + 1 (k+ 1)^3 + 2(k +1) / 3

OpenStudy (irishboy123):

just plough on with it. you have \(\dfrac{ (k+ 1)^3 + 2(k +1) }{ 3}\) \(= \dfrac{ k^3 + 3k^2 + 3k + 1 + 2(k +1) }{ 3}\) \(= \dfrac{k^3 + 2k}{3} + \dfrac{3k^2 + 3k + 3}{3}\) can you se where this is going?!?! just needs finishing off